OpenStudy (anonymous):

find the solutions of the equation which lie in [o,2pi] sec x= -2 ?

terenzreignz (terenzreignz):

Confused about the cosecant? Maybe this will cheer you up :) \[\huge \csc \ \theta=\frac1{\sin \ \theta}\]

OpenStudy (anonymous):

there is sec . made a mistake

OpenStudy (anonymous):

nothing for cheering me up i guess so far:)

terenzreignz (terenzreignz):

Okay... then maybe THIS will cheer you up :) \[\huge \sec \ \theta=\frac1{\cos \ \theta}\]

OpenStudy (anonymous):

yeah but what will be reference angle ?

OpenStudy (anonymous):

cos x = -1\2

terenzreignz (terenzreignz):

^That's right :) And there are only two values of x between 0 and 2pi where that's true :D

OpenStudy (anonymous):

cosx = cosy x = n pie + - y

OpenStudy (anonymous):

yeah you right ! but can u write those values for me :P

OpenStudy (anonymous):

?

OpenStudy (anonymous):

every body giving me links :D not solution of my question :(

terenzreignz (terenzreignz):

But, it's just a matter of looking at that unit circle and pinpointing the cases where cosine equals -1/2

OpenStudy (anonymous):

no usage of reference angle?

terenzreignz (terenzreignz):

The reference angles are marked out in the unit circle :) In degrees and in radiance... for your convenience ;)

OpenStudy (anonymous):

woooo thx got that i just starting to love trignometary :D

OpenStudy (anonymous):

*trigonometry

terenzreignz (terenzreignz):

Awesome :)